OpenStudy (anonymous):

i neen help on 10=2/7n+4 i have to find the value of n and check it

OpenStudy (anonymous):

is it 2/7n + 4 or 2/(7n+4) ??

OpenStudy (anonymous):

10 = \[10=\frac{ 2 }{ 7 }n+4\]

OpenStudy (anonymous):

ignor the first 10 = that I put first thing you do is combine like terms. First off subtract 4 from both sides. 10 = 2/7n + 4 -4 - 4 --------------- 6 = 2/7n Now you have to divide both sides by 2/7 which will equal 2/7 2/7 ---------- 0.04761904761 Now do you have to put that in a fraction form or you have to round to the nearest tenth or hundred

OpenStudy (anonymous):

i have to put it on fraction form and check it too